OpenStudy (anonymous):

1 carpet has aperimeter of 220 inches the length of carpet is 22 inches more than the width.What are dimensions

OpenStudy (mysesshou):

P=220 = 2L+2W L=22+W P = 2(22+W) +2W = 220 solve for W then plug back in again for L

OpenStudy (anonymous):

can you set this up algebraically

OpenStudy (mysesshou):

P = 44 + 2W +2W = 220 44 +4W = 220 4W = 176 W = 44 L = 22 + W = 22 + 44 = 66 L = 66 check: P = 2W + 2L = 2*44 + 2*66 = 88 + 132 = 220

OpenStudy (mysesshou):

is this what you were looking for? makes sense ok?

OpenStudy (mysesshou):

my writing was a continuation of the beginning, for solving

OpenStudy (mysesshou):

dividing by 2 like Luis Rivera did works too. it is simpler. :)